2. (4) Newton’s 2nd law is F = ma. What must be the acceleration of a 4.0 kg box being pulled along a friction-free surface if the net pulling force is 20.0 N? Show all work.
Answer #1

Applied force = 20 N

Mass of box = 4 kg

Using Newton's second law

F = m×a

20 = 4×a

a = 20/4 = 5 m/s​​​​​​2

therefore the acceleration of the 4 kg box when pulled by a force of 20 N = 5 m/s 2